In investing $5,900 of a couple's money, a financial planner put some of it into a savings account paying 5% annual simple interest. The rest was invested in a riskier mini-mall development plan paying 13% annual simple interest. The combined interest earned for the first year was $583. How much money was invested at each rate?
Answer by addingup(3677)

0.05x + 0.13(5,900 - x) = 583
0.05x + 767 - 0.13x = 583
-0.08x = -184
x = 2,300 this is the amount invested at 5%
5,900 - 2,300 = 3,600 this is the amount invested at 13%